OpenStudy (anonymous):

HELP WOULD BE MUCH APPRECIATED! What is the sum of a 7-term geometric series if the first term is -11, the last term is -45,056, and the common ratio is -4? a) -143,231 b) -36,047 c) 144,177 d) 716,144

OpenStudy (anonymous):

danny what are you doing is this homework

OpenStudy (anonymous):

i have like 60 problems and i did most of them exept for these 5 i just posted... i did all i could and im posting these ones on here that i have no idea how to do. sorry. @Jack174

OpenStudy (anonymous):

do you go to school with other kids

OpenStudy (anonymous):

ya this is algabra 2 second semester summer school the course ends tomorrow and im trying to turn in this last assignment by midnight (8 minutes) @Jack174 i have to take my final tomorrow tooo!!!!

OpenStudy (anonymous):

i go to school with other kids... and its online summer school.

OpenStudy (anonymous):

wow that sucks, what grade are you in

OpenStudy (anonymous):

im a junior going to be a senior next year.

OpenStudy (anonymous):

and what were your grades junior year

OpenStudy (anonymous):

lol got any plans for life after high school

OpenStudy (anonymous):

do you think i could get the answer bro??

OpenStudy (anonymous):

is it an online problem?

OpenStudy (anonymous):

whats the family business

OpenStudy (anonymous):

how long do you have to turn in this homework?

OpenStudy (anonymous):

Id help you man, but im really lazy

OpenStudy (anonymous):

aww dang bro.

OpenStudy (kropot72):

Use the formula for the sum of n terms as follows: \[S _{n}=\frac{a(1-r ^{n})}{1-r}\] \[S _{7}=\frac{-11[1-(-4)^{7}]}{1-(-4)}=you\ can\ calculate\]